OpenStudy (anonymous):

what is the max perimeter all all rectangles that can be inscribed in x^2/a^2 + y^2/b^2 = 1

OpenStudy (anonymous):

am trying to type, software glitchy

hartnn (hartnn):

so, we have to maximize 2(x+y) solve for y in x^2/a^2 + y^2/b^2 = 1 then put it back in f(x)=x+.... then 'x' in f'(x)=0

hartnn (hartnn):

u know how to maximize, right ?

OpenStudy (anonymous):

Why maximize 2(x + y)? The perimeter is 2(a + b) isnt it?

hartnn (hartnn):

you'll get x and y in terms of a and b then you put those values in P =2(x+y) and you get perimeter in terms of a and b. and its not 2(a+b)

OpenStudy (anonymous):

oh ya sorry, i swapped the variables on ellipse y = b/a sqrt(a^2 - x^2) y' = -(b x)/(a sqrt(a^2-x^2)) p = 2(x + y) p' = 2(1 + y') p' = 0 whenever y' = -1 y' = -(b x)/(a sqrt(a^2-x^2)) = -1 x = a^2/sqrt(a^2 + b^2) y = b/a sqrt(a^2 - [a^2/sqrt(a^2 + b^2)]^2) x + y = sqrt(a^2+b^2) 2(x + y) = 2 sqrt(a^2 + b^2) but (sorry glitchy google chrome/openstudy)

OpenStudy (anonymous):

but

OpenStudy (anonymous):

2(x+y) only gives the max perimeter of upper portion of ellipse , would it not be 4(x+y)

OpenStudy (anonymous):

i solve for x, but

hartnn (hartnn):

nevermind, got it.

hartnn (hartnn):

you mean to say, that you just took, positive value of 'y', thats why the 'upper' part ?

OpenStudy (anonymous):

ya , for the upper region of the ellipse, i see thank you you are clever

OpenStudy (anonymous):

changing browsers though never use chrome here

hartnn (hartnn):

and since you get the same thing when you do it for negativevalue of 'y', you say,its 2+2=4 ? i don't think thats the case, you take positive value of y because perimeter can't be negative. i still think its 2(x+y) only.

OpenStudy (anonymous):

what do you obtain ? but y' = -1 implies x = +- a^2/(a^2 + b^2) y = +- b/a sqrt(a^2 - x^2) i cannot ignore the sign can i ? perimeter cannot be negative but y values can be. perimeters would be just absolute value of it

hartnn (hartnn):

x=+- .... so if you addd those 'x' values, you get 0 :P and theres no reason to subtract..... i don't have words to explain properly, but it'll be 2(x+y) only.

hartnn (hartnn):